Consider the given vector equation. r(t) = (2t – 5, t2 + 4) (a) Find r(t). r(t)
Answer #1

r(t) = Lat-S, +244) differen hate wit it lt= ($, (242), (42+4)* dt 2 2010 21:07 r(t) = 12, 27 Aug or git at Are = 2 ( 1 + th